# Most discriminative stimuli for functional cell type identification

### 29 Nov 2023 v1

@@ 5 Results
MDS provide time-efficient on-the-fly cell type assignments. As MDS condense the unique response features of each cell type into a single, short stimulus, we next asked if they allow us to assign neurons to types faster than conventional methods. For a fair comparison with conventional approaches, we considered the clusters predicted by the classifier purely trained on the “chirp” responses (see above) as a baseline. For a comparable assessment between clusterings, we matched the 17 “conventional” clusters from the chirp-classifier to the seven MDS clusters, merging types by the Baden et al. (2016) hierarchy into TODO broader classes. The baseline classifier retrained on this task predicted the correct group label with 73.5% accuracy on a held-out test set of neurons. Recording the required chirp responses (four stimulus repetitions) and assigning functional type labels for new neurons requires 2 min 12 s with this baseline classifier.
@@
Marmoset retinal ganglion cells. Next, we demonstrated that our approach is robust across species and recording techniques, using a dataset with the spiking responses of 167 marmoset RGCs to naturalistic videos recorded with multi-electrode arrays. MDS clustering yielded four functional clusters (Fig. 4). Even though the data contained only seven ON cells, the MDS algorithm recovered and even split them into a fast (cluster 4) and a slow (cluster 3) ON cluster. OFF cells were also grouped into a fast and slow cluster (cluster 1 and 2, respectively). These clusters likely correspond to ON and OFF midget (slow) and parasol (fast) cells previously described in the primate retina (Field & Chichilnisky, 2007) and are in good agreement with a clustering baseline based on the spike-triggered averages on white noise stimuli (see Appendix A.3 and Fig. 10).
@@
Macaque visual cortex area V4. Finally, we showed that our approach extends beyond the retina and also works for brain areas with little domain knowledge about cell types. For that, we tested MDS clustering on spiking responses of neurons in macaque area V4, stimulated by gray-scale naturalistic images (Willeke et al., 2023). We discarded clusters containing less than six neurons in the test set, yielding 12 well separated MDS clusters (Fig. 5). Here, the MDS were individual images, which showed complex patterns with curvature or texture, typical for V4 (Bashivan et al., 2019; Willeke et al., 2023). Interpreting these MDS suggests that V4 might compute such complex features on the cell type-level. Repeating clustering with a different seed and ten instead of five initial clusters resulted in the same number of MDS clusters (Appendix Fig. 11) with reasonable similarity (ARI: 0.53). Our results suggest that the MDS algorithm can also be used successfully in computationally more complex brain areas such as primate V4.
@@ 6 Discussion
We presented a novel clustering algorithm that alternates between assigning neurons into functional clusters and optimizing a most discriminative stimulus (MDS) for each functional cluster. We empirically validated our algorithm on three datasets across species and data modalities (gray-scale and multi-channel image and video stimulation, multi-electrode array recordings and calcium-imaging). For mouse retinal ganglion cells (RGCs), we successfully benchmarked the resulting functional clusters against an established functional taxonomy from previous work (Baden et al., 2016) and found a good agreement. While the taxonomy of marmoset RGCs is less well refined than in the mouse, our MDS clustering found clusters that correspond to the four major RGC types.
Data-driven, interpretable cell type discovery with MDS. While previous approaches required domain knowledge to hand-craft fingerprinting stimuli (Farrow & Masland, 2011; Baden et al., 2016; Goetz et al., 2022) and response features for clustering (Farrow & Masland, 2011), our approach identifies functional cell types in a data-driven way by choosing the most informative stimuli automatically. Therefore, our approach is particularly advantageous in situations where domain knowledge is not available and functional cell types are yet to be discovered, such as in later processing stages like primary visual cortex (V1) or V4. Thus, our method could greatly facilitate discovery and interpretation of new functional cell types. To demonstrate this in a setting with little established domain knowledge about functional cell types, we applied our algorithm to cortical area V4, where only first hints to functional groups exist (Willeke et al., 2023). Without incorporating domain knowledge into our model, MDS clustering successfully identified plausible clusters. Inspecting their MDS allowed us to reason about their unique functional properties, as each MDS highlights the visual feature that is uniquely captured by a cluster, and what distinguishes it from the features processed by other types. For instance, the functional RGC type “ON contrast suppressed” has been known since 2016 (Baden et al., 2016), but only recently Hoefling et al. (2022) found that this cell type responds strongly to center color opponency – a property that was not revealed by previous fingerprint stimuli, but is clearly visible in our MDS.
@@
Comparison to other functional clustering methods. Willeke et al. (2023) reported first hints towards functional groups in area V4. They used a digital twin to optimize a maximally exciting image for each single cell and clustered these MEIs in image-space. In contrast, our algorithm optimizes most discriminative stimuli that capture the unique functional properties and distinguish neurons by their responses on the cluster level. To cluster cells by function in primary visual cortex (V1), Ustyuzhaninov et al. (2022) investigated similarities of digital-twin feature weights (associated with a neuron’s function) and perform in-silico experiments designed based on expert knowledge. Their method relies on the (unverified) assumption that such weights actually represent a neuron’s function, and it does not provide interpretable features such as the most discriminative image. In contrast, our algorithm clusters directly based on the functional responses, without the need for expert knowledge or to extract parameters from the digital twin. Unlike both of these approaches, our clustering algorithm is the first that has actually been validated on empirical data where something akin to biological ground truth exists (Baden et al., 2016). In addition, our approach can identify a neuron’s cell type in new experiments quickly and without model training, while both other approaches would need to retrain their digital twin.
Practical advantages. MDS clustering is applicable in many scenarios: It works with any dataset rich enough to train a digital twin, which is the case for many recent experimental studies (Yamins et al., 2014; Bashivan et al., 2019; Lurz et al., 2020; Sinz et al., 2018; Hoefling et al., 2022; Willeke et al., 2022; Turishcheva et al., 2023; Wang et al., 2023). Importantly, once the cell types and their MDS have been established, the MDS alone can be used to quickly identify cell types in new experiments without the digital twin. In contrast to the current practice (record all neurons in an area and identify cell types afterwards), our results suggest that MDS allow for fast online cell typing at recording time. That way, researchers interested in rare subpopulations could identify their cell type of interest with its MDS and target it selectively for data collection and experimental interventions.
Limitations and future work. MDS clustering yields clusters purely based on functional information, whereas previous work additionally included anatomical, and in some cases, transcriptomic features to group neurons of similar type (Baden et al., 2016; Goetz et al., 2022). Our clustering approach required the choice of a digital twin model, and while results in principle can depend on the specific model architecture chosen, our MDS and clustering results were consistent for another digital twin architecture choice (see Sec. A.6 and Fig. 13 for details). While the final judge of the faithfulness of our new technique will be an in-vivo experiment, single cell MEIs generated with the same optimizer and digital twins as we used were verified in-vivo (Hoefling et al., 2022; Willeke et al., 2023) in line with further work on in-vivo verification (Bashivan et al., 2019; Walker et al., 2019; Tong et al., 2023; Franke et al., 2022). This does suggest that our results might well generalize to an in-vivo setting.
@@
Our clustering algorithm can be extended in several ways in the future. For example, ON-OFF RGCs are activated by light increments and light decrements, and only optimizing a single, short MDS might not be sufficient to capture both properties and thus fails to distinguish ON-OFF from pure ON or pure OFF cells. However, our method can be extended to optimize multiple MDS per cluster, and pilot results (see Sec. A.5 and Fig. 12) showed this is a promising direction to solve this issue. Similarly, “OFF suppressed 2” cells exhibit a high baseline firing rate that is mostly suppressed by stimulation. For such cell types, searching for a maximally suppressing stimulus might be another promising direction. Another promising direction for future work is to test and validate our method on new functional recordings and transcriptomic data, e.g. by testing experimentally which cells are driven by our MDS and later identifying their transcriptomic types by PatchSeq (Liu et al., 2020) or spatial transcriptomics (Alon et al., 2021) in the same tissue. As currently set up, MDS clustering assumes discrete cell types. Extending it towards capturing recently reported continuity within or between cell types (Tasic et al., 2016; Harris et al., 2018; Tasic et al., 2018; Stanley et al., 2020; Scala et al., 2021) is another interesting direction.
@@
Conclusions. Overall, we think that our general-purpose functional cell type clustering algorithm and clusters’ most discriminative stimuli will be a useful tool for the neuroscience community. We expect that this approach will spark further work on exploring functional groups across the visual system – especially in areas where fingerprinting stimuli are to date unavailable – and help designing experiments previously limited by experiment time.
@@
We will make our code, seeds, and used Python environment publicly available for reproducibility at publication time. Furthermore, we provide detailed method descriptions and parameters in sections Most discriminative stimulus clustering algorithm, Data and implementation details, and Appendix. For all results we quantified reproducibility of MDS clustering across runs with varying parameter initializations and we provide according visualizations in Figs. 6, 9 and 11.
@@ Acknowledgments
The authors would like to thank (in alphabetic order): Catie Nealley, Emmanouil Froudarakis, Federico D’Agostino, Gabrielle Rodriguez, Jiakun Fu, Katrin Franke, Kayla Ponder, Polina Turishcheva, Saumil Patel, Taliah Muhammad, Tori Shinn. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting MFB, KFW, and SM. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant agreement No. 101041669). This publication was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 432680300 - SFB 1456. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP4, project number: 276693517. MB and PB are members of the Machine Learning Cluster of Excellence “Machine Learning — New Perspectives for Science”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. We thank the European Union (ERC, “NextMechMod”, ref. 101039115). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. We also thank the Hertie Foundation for support. This work was supported by the Tübingen AI Center. The work was supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DoI/IBC) contract number D16PC00003. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. AST acknowledges support from NSF NeuroNex grant 1707400. AST acknowledges support from National Institute of Mental Health and National Institute of Neurological Disorders And Stroke under Award Number U19MH114830 and National Eye Institute award numbers R01 EY026927 and Core Grant for Vision Research T32-EY-002520-37. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DoI/IBC, or the U.S. Government. The authors gratefully acknowledge the computing time made available to them on the high-performance computers HLRN-IV at GWDG at the NHR Center NHR@Göttingen.
@@ A.2 Mouse Retina details
Figure 7: Results of a typical clustering run (run with median ARI compared to results in Fig. 2). Results on held-out test neurons and empirical comparison to Baden et al. (2016) types shown. A. Cluster averaged digital twin response to the optimized MDS. Elements within a column normalized to highest value. B. Confusion matrix between neuron’s MDS cluster assignment in Fig. 2 and the most typical comparison run. C. Confusion matrix between MDS clusters and Baden et al. (2016) types. Elements within rows (across Baden et al. (2016) types) were normalized to sum to 1, annotated numbers display the number of neurons in percent.
@@
Figure 8: Mean traces of held-out test neurons in response to the most discriminative stimuli (left) grouped by the Baden et al. (2016) RGC type. Group MEIs for each cluster are visualized on top of the plot. The extend of the x and y axis is shared across all plots with response traces.
@@
Digital twin training. Each member was initialized with a different seed and trained independently using Adam (Kingma & Ba, 2015) minimizing a Poisson loss:
$\sum_{n=1}^{N}\left(\bm{\hat{r}^{(n)}}-\bm{r^{(n)}}\log\bm{\hat{r}^{(n)}}\right)$
@@ A.3 Marmoset Retina Details
Robustness check. As a simple robustness check of MDS clustering, we compared to a typical comparison run (the one with mean ARI compared to Fig. 4 selected across nine comparison runs varying in the number of initial clusters, intial cluster assignment, and MDS initialization. The resulting clusters and average responses are visualized in Fig. 9.
@@
Digital twin training. We trained the CNN using Adam optimizer (Kingma & Ba, 2015) optimizing Poisson loss Sec. A.2 on 16 out of 20 available 5 minute trials. The batch size was 16 and used early stopping on the correlation between predicted and recorded responses on the validation set which consisted of the 4 remaining trials. We reduced learning rate by a factor of 0.1 each time the validation correlation did not improve for 5 epochs. The initial learning rate was 0.008 and the minimal learning rate we allowed was $8\cdot 10^{-5}$. We stopped training early if the validation correlation did not improve for 50 epochs.
MDS optimization. To optimize the MDS on this dataset we used an ensemble of 5 of the above described models varying in their initialization before model training. We also tailored a few of the parameters of MDS optimization to this dataset, optimizing 39 frames ×$\times\,(80\times 90)$ pixels one-channel MDS videos (85 Hz frame-rate) with a learning rate of 4 and chose ten initial clusters. Unlike for the mouse retina data, here we fed the predicted activity of only the last time-bin to the discriminative objective function (Eq. 1). The MDS input was constraint with an L2-norm of 3.5 and the pixel values were clipped to stay within the range of -1 and 1. We optimized the stimulus for a maximum of 10,000 optimization steps, and for 10 steps during splitting, or until an exponential moving average of the objective (Eq. 1) did not improve for 100 optimization steps. During splitting, we created two new sub-cluster candidates and terminated splitting after 100 iterations. As neuron’s responses are not normalized in this data, we post-hoc normalized digital twin predictions such that for each neuron, predictions were standard-normal distributed across the digital twin’s training stimuli, ensuring single cell responses do not dominate the within-cluster mean.
@@
STA-based cell clustering. To be able to compare our MDS clustering of marmoset RGCs to an alternative clustering approach, we determined each cell’s RF size as the area within the 1.5-sigma contour of a Gaussian fitted to the cell’s spatial component of the spike triggered average (STA) measured under spatio-temporal white-noise stimulation. The STA’s temporal component and the RF size were taken as a feature vector for each cell. The feature vectors were normalized and their dimensionality was reduced using Principal Component Analysis (PCA). We selected as many principal components as necessary to explain 90% of the variance. We then used KMeans++ (Arthur & Vassilvitskii, 2007) on the reduced-dimensionality feature vectors to determine 4-6 initial clusters. These clusters were then manually curated to ensure each cluster had similarly-sized cells and tiled the retinal surface, resulting in a total of nine clusters. However, the taxonomy of functional types in the marmoset is less understood compared to the mouse and a concise interpretation of these clusters is not yet established, so they cannot be treated as ground truth. Comparing these clusters to our MDS approach yielded overall good agreement (Fig. 10).
Data recording. All marmoset data were obtained by recording RGC spikes extracellularly from isolated retina placed on a MEA, as explained in (we will add this reference in the camera-ready version to keep authors anonymous for double-blind review). Marmoset retinal tissue was obtained immediately after euthanasia from animals used by other researchers, in accordance with national and institutional guidelines and as approved by the institutional animal care committee of the German Primate Center and by the responsible regional government office (Niedersächsisches Landesamt für Verbraucherschutz und Lebensmittelsicherheit, permit number 33.19-42502-04-20/3458).

### 14 Mar 2024 v2

@@ 5 Results
MDS provide time-efficient on-the-fly cell type assignments. As MDS condense the unique response features of each cell type into a single, short stimulus, we next asked if they allow us to assign neurons to types faster than conventional methods. For a fair comparison with conventional approaches, we considered the clusters predicted by the classifier purely trained on the “chirp” responses (see above) as a baseline. For a comparable assessment between clusterings, we merged the 17 “conventional” clusters from the chirp-classifier according to the Baden et al. (2016) hierarchy and matched them to the seven MDS clusters. The baseline classifier retrained on this task predicted the correct cluster label with 73.5% accuracy on a held-out test set of neurons. Recording the required chirp responses (four stimulus repetitions) and assigning functional type labels for new neurons requires 2 min 12 s with this baseline classifier.
@@
Marmoset retinal ganglion cells. Next, we demonstrated that our approach is robust across species and recording techniques, using a dataset with the spiking responses of 167 marmoset RGCs to naturalistic videos recorded with multi-electrode arrays. MDS clustering yielded four functional clusters (Fig. 4). Even though the data contained only seven ON cells, the MDS algorithm recovered and even split them into a slow and fast ON cluster (cluster 3 and 4, respectively). OFF cells were also grouped into a fast and slow cluster (cluster 1 and 2, respectively). These clusters likely correspond to ON and OFF midget (slow) and parasol (fast) cells previously described in the primate retina (Field & Chichilnisky, 2007) and are in good agreement with a clustering baseline based on the spike-triggered averages on white noise stimuli (see Appendix A.3 and Fig. 10).
@@
Macaque visual cortex area V4. Finally, we showed that our approach extends beyond the retina and also works for brain areas with little domain knowledge about cell types. For that, we tested MDS clustering on spiking responses of neurons in macaque area V4, stimulated by gray-scale images (Willeke et al., 2023). We discarded clusters containing less than six neurons in the test set, yielding 12 well separated MDS clusters (Fig. 5). Here, the MDS were individual images, which showed complex patterns with curvature or texture, typical for V4 (Bashivan et al., 2019; Willeke et al., 2023). Interpreting these MDS suggests that V4 might compute such complex features on the cell type-level. Repeating clustering with a different seed and ten instead of five initial clusters resulted in the same number of MDS clusters (Appendix Fig. 11) with reasonable similarity (ARI: 0.53). Our results suggest that the MDS algorithm can also be successful in computationally more complex brain areas such as primate V4.
@@ 6 Discussion
We presented a novel clustering algorithm that alternates between assigning neurons into functional clusters and optimizing a most discriminative stimulus (MDS) for each functional cluster. We empirically validated our algorithm on three datasets across species, visual processing stages (mouse/marmoset RGCs, macaque V4) and data modalities (gray-scale and multi-channel image and video stimulation, multi-electrode array recordings and calcium-imaging), and showed that MDS allow to identify cell types faster than traditional methods.
@@
Comparison to other functional clustering methods. Willeke et al. (2023) reported hints towards functional groups in macaque V4. They optimized a maximally exciting image for each single cell and clustered these MEIs. In contrast, our algorithm computes MDS capturing unique functional properties that distinguish clusters of neurons based on their responses. In mouse V1, Ustyuzhaninov et al. (2022) clustered non-interpretable digital-twin model parameters they assume to represent a neuron’s function. In contrast, our algorithm clusters directly based on responses to interpretable MDS, not requiring expert knowledge to extract parameters from the digital twin. Unlike both of these approaches, we empirically validated our clustering algorithm on data where type labels akin to biological ground truth exist (Baden et al., 2016). In addition, once the cell types and their MDS have been established, our approach can identify a neuron’s type in new experiments without the digital twin, while both other approaches would need to retrain the digital twin. This suggests that MDS allow for fast online cell typing at recording time, enabling researchers to target a specific cell type they are interested in for data collection and experimental interventions.
Data-driven and interpretable cell type discovery with MDS. While previous clustering approaches required domain knowledge to hand-craft fingerprinting stimuli (Farrow & Masland, 2011; Baden et al., 2016; Goetz et al., 2022), our approach identifies functional cell types by automatically generating a fingerprinting stimulus: the MDS. Therefore, our approach is particularly advantageous when domain knowledge is scarce and functional types are yet to be discovered, for example in V1 or – as we demonstrated – in V4: here, without incorporating domain knowledge, our clustering approach successfully identified plausible clusters.
The resulting MDS facilitate interpretability of a cluster’s function, as it highlights a cluster’s unique visual feature that distinguishes it from the features processed by other groups. For instance, the mouse RGC type “ON contrast suppressed” has been a subject of study for some time (Tien et al., 2015; Baden et al., 2016; Mani & Schwartz, 2017; Tien et al., 2022), but only recently Hoefling et al. (2022) found that this cell type responds strongly to center color opponency – a property that was not revealed by previous fingerprint stimuli, but is clearly visible in our MDS. For complex visual features (e.g. in V4, Fig. 5), describing a cluster’s function in words might be more difficult and provides an interesting avenue for future work.
Limitations and future work. MDS clustering is applicable for any dataset rich enough to train a digital twin, which is the case for many recent experimental studies (Yamins et al., 2014; Bashivan et al., 2019; Lurz et al., 2020; Sinz et al., 2018; Hoefling et al., 2022; Willeke et al., 2022; Turishcheva et al., 2023; Wang et al., 2023). While a systematic investigation of how the choice of a specific digital twin model affects downstream results is to date unavailable, our MDS and clusters were consistent with another digital twin architecture choice (Sec. A.6 and Fig. 13), suggesting a certain degree of robustness to model details. Another promising direction for future work is to apply our method to new functional recordings where no comparison clustering exists and to validate it with transcriptomic data, e.g. by testing experimentally which cells are driven by our MDS and subsequently identifying their transcriptomic types by PatchSeq (Liu et al., 2020) or spatial transcriptomics (Alon et al., 2021) in the same tissue. While the final judge of the MDS’ faithfulness would be an in vivo experiment, single cell MEIs generated with the same optimizer and digital twins as we used were verified in vivo (Willeke et al., 2023) and ex vivo (Hoefling et al., 2022), suggesting that our results might well generalize to an in vivo setting – similar to other work on in vivo verification (Bashivan et al., 2019; Walker et al., 2019; Franke et al., 2022; Tong et al., 2023).
@@
Algorithm extensions. ON-OFF RGCs are stimulated by both light increases and decreases. Solely optimizing a single, short MDS may not effectively capture both properties and thus fails to distinguish ON-OFF from pure ON or pure OFF cells. To address this, our method can be expanded to optimize multiple MDS for each cluster, showing promising initial results (see Sec. A.5 and Fig. 12). Additionally, identifying “OFF suppressed 2” cells, characterized by their high baseline firing rate reduced through stimulation, might be improved by searching for a maximally suppressing stimulus. As currently set up, MDS clustering assumes discrete cell types, and future work could extend it towards accounting for recently reported continuity within or between cell types (Tasic et al., 2016; Harris et al., 2018; Tasic et al., 2018; Stanley et al., 2020; Scala et al., 2021).
@@
Conclusions. Our general-purpose functional cell type clustering algorithm and clusters’ most discriminative stimuli could be a useful tool for the neuroscience community and spark further work on exploring functional groups across the visual system – especially in areas where no fingerprinting stimuli are available – and help designing experiments previously limited by experiment time.
@@
We will make our code, seeds, and used Python environment publicly available for reproducibility at https://github.com/ecker-lab/most-discriminative-stimuli. Furthermore, we provide detailed method descriptions and parameters in sections Most discriminative stimulus clustering algorithm, Data and implementation details, and Appendix. For all results we quantified reproducibility of MDS clustering across runs with varying parameter initializations and we provide according visualizations in Figs. 6, 9 and 11.
@@ Acknowledgments
The authors would like to thank (in alphabetic order): Catie Nealley, Emmanouil Froudarakis, Federico D’Agostino, Gabrielle Rodriguez, Jiakun Fu, Katrin Franke, Kayla Ponder, Polina Turishcheva, Saumil Patel, Taliah Muhammad, Tori Shinn. MFB, KFW, and SM thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant agreement No. 101041669). This publication was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 432680300 - SFB 1456. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP4, project number: 276693517. MB and PB are members of the Machine Learning Cluster of Excellence “Machine Learning — New Perspectives for Science”, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. We thank the European Union (ERC, “NextMechMod”, ref. 101039115). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. We also thank the Hertie Foundation for support. This work was supported by the Tübingen AI Center. The work was supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DoI/IBC) contract number D16PC00003. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. AST acknowledges support from NSF NeuroNex grant 1707400. AST acknowledges support from National Institute of Mental Health and National Institute of Neurological Disorders And Stroke under Award Number U19MH114830 and National Eye Institute award numbers R01 EY026927 and Core Grant for Vision Research T32-EY-002520-37. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DoI/IBC, or the U.S. Government. The authors gratefully acknowledge the computing time made available to them on the high-performance computers HLRN-IV at GWDG at the NHR Center NHR@Göttingen.
@@ A.2 Mouse Retina details
Figure 7: Results of a typical clustering run (run with median ARI compared to results in Fig. 2). Results on held-out test neurons and empirical comparison to Baden et al. (2016) types shown. A. Cluster averaged digital twin response to the optimized MDS. Elements within a column normalized to highest value. B. Confusion matrix between neuron’s MDS cluster assignment in Fig. 2 and the most typical comparison run. C. Confusion matrix between MDS clusters and Baden et al. (2016) types. Elements within rows were normalized to sum to 1, annotated numbers display the number of neurons in percent.
@@
Figure 8: Mean traces of held-out test neurons in response to the most discriminative stimuli (left) grouped by the Baden et al. (2016) RGC type. Group MEIs for each cluster are visualized on top of the plot. The extend of the x- and y-axis is shared across all plots with response traces.
@@
Digital twin training. Each member was initialized with a different seed and trained independently using the Adam optimizer (Kingma & Ba, 2015) minimizing a Poisson loss,
$\mathcal{L}_{\text{Poisson}}=\sum_{n=1}^{N}\left(\bm{\hat{r}^{(n)}}-\bm{r^{(n)}}\ln\bm{\hat{r}^{(n)}}\right)\ ,$ (2)
@@ A.3 Marmoset Retina Details
Robustness check. As a simple robustness check of MDS clustering, we compared to a typical comparison run (the one with mean ARI compared to Fig. 4 selected across nine comparison runs varying in the number of initial clusters, intial cluster assignment, and MDS initialization). The resulting clusters and average responses are visualized in Fig. 9.
@@
Digital twin training. We trained the CNN using the Adam optimizer (Kingma & Ba, 2015) optimizing the Poisson loss (Eq. 2) on 16 out of 20 available 5 minute trials. The batch size was 16 and we used early stopping on the correlation between predicted and recorded responses on the validation set which consisted of the 4 remaining trials. We reduced the learning rate by a factor of 0.1 each time the validation correlation did not improve for 5 epochs. The initial learning rate was 0.008 and the minimal learning rate we allowed was $8\cdot 10^{-5}$. We stopped training early if the validation correlation did not improve for 50 epochs.
MDS optimization. To optimize the MDS on this dataset we used an ensemble of 5 of the above described models varying in their initialization before model training. We also tailored a few of the parameters of MDS optimization to this dataset, optimizing 39 frames ×$\times\,(80\times 90)$ pixels one-channel MDS videos (85 Hz frame-rate) with a learning rate of 4 and chose ten initial clusters. Unlike for the mouse retina data, here we fed the predicted activity of only the last time-bin to the discriminative objective function (Eq. 1). The MDS input was constraint with an $L_{2}$-norm of 3.5 and the pixel values were clipped to stay within the range of -1 and 1. We optimized the stimulus for a maximum of 10,000 optimization steps, and for 10 steps during splitting, or until an exponential moving average of the objective (Eq. 1) did not improve for 100 optimization steps. During splitting, we created two new sub-cluster candidates and terminated splitting after 100 iterations. As neuron’s responses are not normalized in this data, we post-hoc normalized digital twin predictions such that for each neuron, predictions were standard-normal distributed across the digital twin’s training stimuli, ensuring single cell responses do not dominate the within-cluster mean.
@@
STA-based cell clustering. To be able to compare our MDS clustering of marmoset RGCs to an alternative clustering approach, we determined each cell’s RF size as the area within the 1.5 standard deviation contour of a Gaussian fitted to the cell’s spatial component of the spike triggered average (STA) measured under spatio-temporal white-noise stimulation. The STA’s temporal component and the RF size were taken as a feature vector for each cell. The feature vectors were normalized and their dimensionality was reduced using Principal Component Analysis (PCA). We selected as many principal components as necessary to explain 90% of the variance. We then used KMeans++ (Arthur & Vassilvitskii, 2007) on the reduced-dimensionality feature vectors to determine 4-6 initial clusters. These clusters were then manually curated to ensure each cluster had similarly-sized cells and tiled the retinal surface, resulting in a total of nine clusters. However, the taxonomy of functional types in the marmoset is less understood compared to the mouse and a concise interpretation of these clusters is not yet established, so they cannot be treated as ground truth. Comparing these clusters to our MDS approach yielded overall good agreement (Fig. 10).
Data recording. All marmoset data were obtained by recording RGC spikes extracellularly from isolated retina placed on a MEA, as explained in Krüppel et al. (2023). Marmoset retinal tissue was obtained immediately after euthanasia from animals used by other researchers, in accordance with national and institutional guidelines and as approved by the institutional animal care committee of the German Primate Center and by the responsible regional government office (Niedersächsisches Landesamt für Verbraucherschutz und Lebensmittelsicherheit, permit number 33.19-42502-04-20/3458).